Given the points (x1,y1)=(-4,8) & (x2,y2)=(6,2) Find m (slope) and b (y-axis intercept) of the equation y=mx+ b
The slope m= and b=

Answer by checkley71(8403)

SLOPE=(2-8)/(6+4) OR SLOPE=-6/10 OR SLOPE=-3/5 AND THE Y INTERCEPT IS
(3/5=X/6) OR 5X=18 OR X=3.6 OR THE Y INTERCEPT IS 3.6+2=5.6 THUS THE EQUATION IS Y=-3/5X+5.6